# https://vcai.mpi-inf.mpg.de/projects/MetaCap/

### v1

@@ 3.3 Occlusion Handling and Fine-tuning
Figure 3: Occlusion Handling. (a) Monocular RGB input. (b) Naive fine-tuned results. (c) Visibility map and proxy image in canonical space. (d) Fine-tuned results with occlusion handling. Our occlusion handling helps regularize largely occluded regions.
[…]
Occlusion Handling. For extreme occlusion cases as it is common in the monocular setting (see also Fig. 3), we further propose a dedicated occlusion handling strategy by leveraging the posed human template. First, we pre-build an implicit human field on one frame of dense images $\mathcal{I}_{f}$ or multiple frames of (in-the-wild) sparse images ^$\hat{\mathcal{I}}_{f}$, and then render virtual views of the canonical space as proxy RGB images $\mathcal{I}^{proxy}$. Next, during the fine-tuning stage, we determine per-vertex visibility by checking whether the vertex of the template is visible in any of the sparse input views. Finally, in addition to sampling rays from the input images, we sample additional virtual rays for the occluded regions from $\mathcal{I}^{proxy}$. In practice, we found that the occlusion handling is only required in the case of monocular reconstruction (also see ablations).
@@ 4 Results
Implementation Details. For meta-learning, we perform $M=24$ inner-loop gradient steps with $l^{out}=1.0$ and $l^{in}=1e-4$. The inner loop optimizer is Adam [33] while the outer one uses stochastic gradient descent. Following NeuS [76], we perform a warmup in the inner and outer loop at the beginning of training. If not stated otherwise, we use DDC [22] for space canonicalization, $\sim$80 cameras for meta learning, 4 cameras for fine-tuning in comparisons and ablations. For more details, we refer to the supplemental material.
@@
Metrics and Ground-Truth. To obtain the GT geometry for the dome captures, we leverage the recent state-of-the-art implicit reconstruction method InstantNSR [93] trained on the dense RGB views. To evaluate the rendering quality, we report the peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learned perceptual image patch similarity (LPIPS) [91]. To evaluate geometry results, we report Chamfer distance (CD), point-to-mesh distance (P2S), intersection over union (IOU), cosine normal consistency (NC-Cos), and L2 normal consistency (NC-L2). For all metrics, we sample every 100th frame from the test set, where the actor is performing unseen motions, and report the average. We provide an illustration of input cameras and evaluation cameras distribution in the supplementary material.
@@ 4.1 Comparisons
Figure 5: Qualitative Ablation. Here, we show qualitative results for different weight initialization strategies and space canonicalization types. Note that our meta-learning paradigm together with the proposed space canonicalization achieves the best result. Moreover, we highlight that this applies irrespective of the choice of the character model, i.e. SMPL [40] or DDC [22].
@@ 4.2 Ablation Studies
Figure 8: Qualitative In-the-wild Comparison. To evaluate the robustness to in-the-wild conditions, we compare our work and ARAH on WildDynaCap. Our approach achieves significantly better results in terms of appearance and geometric details. (a, d) donate ground truth images at novel view. (b, e) are the results of our method. (c, f) are the results of ARAH.
@@ 4.3 Evaluation on In-the-wild Sequences
5 Limitations
Although our method achieves high-fidelity personalized human reconstruction and rendering from sparse observation, it still has a few limitations. First, our method can be sensitive to template fitting or motion capture results. Second, we do not take the temporal information into consideration. Integrating reliable observations or constraints from adjacent frames like Newcombe et al. [46] may lead to more robust geometry and rendering results. Third, our method can not handle hands very well. Modeling hand with a more fine-grained template and motion capture could be a potential solution.
@@ 6 Conclusion
Acknowledgements: This research was supported by the ERC Consolidator Grant 4DRepLy (770784).

### v2

@@ 3.3 Occlusion Handling and Fine-tuning
Figure 3: Occlusion Handling. Our occlusion handling (abbreviated as OH) helps to regularize largely occluded regions.
At inference, when only the calibrated monocular or sparse-view RGB images $\mathcal{{I}}_{f}$, masks $\mathcal{{O}}_{f}$, and skeletal motion $\mathcal{{M}}_{f,W}$ are available, we initialize our implicit human field with the meta-learned weights $\phi^{*}$, conduct the space canonicalization with the fitted template $\mathbf{X}$ (Sec.3.2), and fine-tune using the loss from Eq.2. This can also be denoted as the task {,$\{\mathcal{L},\{\mathbf{R}_{f}\sim{\mathcal{I}}_{f},{\mathcal{O}}_{f}\},f_{\phi}|{\phi_{0}}=\phi^{*}\}_{f}$.
Occlusion Handling. For extreme occlusion cases, which may occur in the monocular setting (see also Fig. 3), we further propose a dedicated occlusion handling strategy by leveraging the posed human template. Inspired by [77], we first pre-build an implicit human field on one frame of dense images $\mathcal{I}_{f}$ or multiple frames of (in-the-wild) sparse images ${\mathcal{I}}_{f}$, and then render virtual views of the canonical space as proxy RGB images $\mathcal{I}^{\mathrm{proxy}}$. Next, during the fine-tuning stage, we determine per-vertex visibility by checking whether the vertex of the template is visible in any of the sparse input views. Finally, in addition to sampling rays from the input images, we sample additional virtual rays for the occluded regions from $\mathcal{I}^{\mathrm{proxy}}$. In practice, we found that the occlusion handling is only required in the case of monocular reconstruction (also see ablations).
@@ 4 Results
Implementation Details. For meta-learning, we perform $M=24$ inner-loop gradient steps with $l^{out}=1.0$ and $l^{in}=1e-4$. The inner loop optimizer is Adam [34] while the outer one uses stochastic gradient descent. Following NeuS [82], we perform a warmup in the inner loop at the beginning of training. If not stated otherwise, we use DDC [22] for space canonicalization, $\sim$80 cameras for meta learning, 4 cameras for fine-tuning in comparisons and ablations. For more details, please refer to the supplemental material.
@@
Metrics and Ground-Truth. To obtain the ground truth geometry for the dome captures, we leverage the recent state-of-the-art implicit reconstruction method InstantNSR [99] trained on the dense RGB views. To evaluate the rendering quality, we report the peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and learned perceptual image patch similarity (LPIPS) [97]. To evaluate geometry results, we report Chamfer distance (CD), point-to-mesh distance (P2S), intersection over union (IOU), cosine normal consistency (NC-Cos), and L2 normal consistency (NC-L2). For all metrics, we sample every 100th frame from the test set, where the actor is performing unseen motions, and report the average. We provide an illustration of input cameras and evaluation cameras distribution in the supplementary material.
@@ 4.1 Comparisons
Figure 5: Qualitative Ablation. We show results for different weight initialization strategies and space canonicalization types. Our meta-learning paradigm, together with the proposed space canonicalization, achieves the best result. We highlight that this applies irrespective of the choice of the character model (SMPL [43] or DDC [22]).
@@ 4.2 Ablation Studies
Figure 8: Qualitative In-the-wild Comparison. To evaluate the robustness to in-the-wild conditions, we compare our work and ARAH on WildDynaCap. Our approach achieves significantly better results in terms of appearance and geometric details.
@@ 5 Conclusion
Limitations. Although our method achieves high-fidelity personalized human reconstruction and rendering from sparse observation, it still has a few limitations. First, our method can be sensitive to template fitting or motion capture results. Second, we do not take the temporal information into consideration. Integrating reliable observations or constraints from adjacent frames like Newcombe et al. [50] may lead to more robust geometry and rendering results. Third, our method can not handle hands very well. Modeling hand with a more fine-grained template and motion capture could be a potential solution.
Acknowledgements
This research was supported by the ERC Consolidator Grant 4DRepLy (770784).